Please help me finish solving this triangle if A = 150°, b = 4.8, a = 9.4
sin(150)/9.4 = sin(b)/4.8

Answer by Alan3354(69443)

Please help me finish solving this triangle if A = 150°, b = 4.8, a = 9.4
sin(150)/9.4 = sin(b)/4.8
Use CAPS for the angles.
sin(150)/9.4 = sin(B)/4.8
Find sin(B)
sin(B) = 4.8*sin(150)/9.4 =~ 0.255319
B =14.7925 degs
-----
C = 180 -(A+B) = 15.2075 degs
===================
c/sin(C) = a/sin(A)
Solve for side c.